# Training Robust Graph Neural Networks by Modeling Noise Dependencies

### v1

@@ B.1 Details of regularizing the inference of $Z_{A}$
We aim to regularize $Z_{A}$ to align with our prior knowledge by minimizing the KL divergence between the latent graph structure $\hat{\mathbf{A}}$ and prior graph structure $\mathbf{A}^{\text{p}}=\{{a}^{\text{p}}_{ij}\}_{N\times N}$, where ${a}^{\text{p}}_{ij}$ is sampled from a Bernoulli distribution with a probability as given by the $\gamma$-hop subgraph similarity.
@@ Appendix C Further Discussion on DANG
C.1 Additional Real-world Examples of DANG
@@
  • Biological Networks: In addition to the user-item graphs, DANG manifests in the domain of single-cell RNA-sequencing (scRNA-seq). Specifically, in this graph the primary resource is a cell-gene count matrix. A cell-cell graph is commonly employed for downstream tasks, where each cell and its corresponding gene expression are represented as a node and node feature, respectively, and the cell type is considered a node label. However, the node feature, representing gene expression derived from the cell-gene count matrix, often contains noise due to various reasons, such as the dropout phenomenon (Hicks et al., 2018) and batch effect (Shaham et al., 2017). Since the cell-gene count matrix is the main resource for generating the cell-cell graph (Wang et al., 2021; Xiong et al., 2023; Yun et al., 2023), such noise acts as a significant obstacle in designing an effective graph structure. Additionally, cell types are annotated using transcripted marker genes, which serve as distinctive features characterizing specific cell types. Noisy node features, therefore, can lead to the misprediction of cell types (node labels). This issue of noise in node features in the biological domain underscores the critical challenge in real-world scenarios.
C.2 Extension of DANG
@@ Appendix D Further Discussion on Da-Gnn
In this section, to further highlight the technical distinction of our work, we compare Da-Gnn with the recent works handling the data noise using variational inference technique.
D.1 Comparison to CausalNL (Yao et al., 2021)
While the implementation of Da-Gnn draws inspiration from the spirit of VAE (Kingma & Welling, 2013) and CausalNL (Yao et al., 2021), we address complex and unique challenges absent in (Kingma & Welling, 2013; Yao et al., 2021). Specifically, the incorporation of $A$ necessitates handling supplementary latent variables and causal relationships, such as $Z_{A}$, $\epsilon_{A}$, $A$$\epsilon_{A}$, $A$$X$, $Y$$A$, $A$$Z_{A}$, each posing non-trivial obstacles beyond their straightforward extension.
  • While (Yao et al., 2021) assumes that $\epsilon$ only causes $X$, DANG posits that $\epsilon$ also causes $A$, denoted as $A$$\epsilon_{A}$. Consequently, DANG requires a novel inference/regularization approach for $\epsilon_{A}$, which is not addressed in (Yao et al., 2021), presenting a distinctive technical challenge.
  • A simplistic uniform prior is employed to regularize the modeling of $Z_{Y}$ in (Yao et al., 2021). However, upon close examination of the relationship $Y$$A$, we advocate for a novel regularization approach for $Z_{Y}$ based on the principle of homophily. This method cannot be elicited through a straightforward application of (Yao et al., 2021) to the graph.
  • By incorporating $A$, $Z_{A}$, and their associated casualties, we address distinct technical challenges, specifically the inference/regularization of $Z_{A}$ and the generation of $A$, which cannot be accommodated by a mere extension of (Yao et al., 2021) to the graph. In particular, we utilize graph structure learning to model $Z_{A}$, and frame the generation of $A$ as an edge prediction task, incorporating novel regularization techniques for both edge prediction and label. Moreover, we regularize $Z_{A}$ leveraging our novel prior knowledge to enhance the accuracy and scalability of inference.
We argue that these components are non-trivial to handle through a straightforward application of (Yao et al., 2021) to the graph domain.
D.2 Comparision to GraphDE (Li et al., 2022b)
GraphDE (Li et al., 2022b) aims to handle situations where out-of-distribution (OOD) samples are present in training data. Given that a noisy sample can be viewed as an OOD sample, GraphDE shares a similar objective with our work. Additionally, GraphDE employs a generative model based on variational inference, relevant to learning approach. However, there are significant distinctions in the DGP that we assume compared to GraphDE. Specifically, GraphDE assumes that a latent variable $e$ determines whether an instance is an OOD sample or not, overlooking the way that the OOD sample is generated. Conversely, the DGP of our proposed DANG characterizes the process by which noisy samples (i.e., OOD sample) are generated (e.g., in a feature-dependent manner). This crucial difference enables a more fine-grained model learning than GraphDE, which may weaken the applicability of GraphDE to complex real-world noise scenarios, such as DANG.
D.3 Comparison to RGIB (Zhou et al., 2023)
RGIB (Zhou et al., 2023) aims at denoising graph structure, primarily centered on the link prediction task. In their work, while $Z_{A}$ and $A$ share the same meaning as in our paper, $Y$ denotes ”edge labels” rather than node labels (i.e., it serves as a binary indicator for the presence of query edges), and $Z_{Y}$ refers to ”clean edge labels” rather than clean latent node labels. These distinctions significantly differentiate their approach from ours. Furthermore, it’s important to note that RGIB primarily addressed noisy graph structures while assuming that node features and node labels are noise-free. In contrast, our proposed method, Da-Gnn, is developed under the more realistic DANG assumption, where node features, graph structures, and node labels all simultaneously contain noise. This fundamental difference underscores the enhanced applicability and robustness of Da-Gnn compared to RGIB.
@@ E.1 Datasets
We evaluate Da-Gnn and baselines on five existing datasets (i.e., Cora (Yang et al., 2016), Citeseer (Yang et al., 2016), Amazon Photo and Computers (Shchur et al., 2018)), and ogbn-arxiv (Hu et al., 2020) and two newly introduced datasets (i.e., Amazon Auto and Amazon Garden) that are proposed in this work based on Amazon review data (He & McAuley, 2016; McAuley et al., 2015) to mimic DANG caused by malicious fraudsters on e-commerce systems (Refer to Appendix E.2.2 for details). The statistics of the datasets are given in Table 4. These seven datasets can be found in these URLs:
@@ E.2.2 Real-world DANG
We have introduced and released two new graph benchmark datasets, i.e., Auto and Garden, that simulate real-world DANG scenarios on e-commerce systems. To construct these graphs, we utilized metadata and product review data from two categories, ”Automotives” and ”Patio, Lawn and Garden,” obtained from Amazon product review data sources (He & McAuley, 2016; McAuley et al., 2015). Specifically, we generated a clean product-product graph where node features are represented using a bag-of-words technique applied to product reviews. The edges indicate co-purchase relationships between products that have been purchased by the same user, and the node labels correspond to product categories. We perform both node classification and link prediction tasks, which are equivalent to categorizing products and predicting co-purchase relationships, respectively.
@@ E.3 Baselines
We compare Da-Gnn with a wide range of noise-robust GNN methods, which includes feature noise-robust GNNs (i.e., AirGNN (Liu et al., 2021)), structure-noise robust GNNs (i.e., ProGNN (Jin et al., 2020), RSGNN (Dai et al., 2022), STABLE (Li et al., 2022a) and EvenNet (Lei et al., 2022)), label noise-robust GNNs (i.e., NRGNN (Dai et al., 2021) and RTGNN (Qian et al., 2023)), and multifaceted noise-robust GNNs (i.e., SG-GSR (In et al., 2024). We also consider WSGNN (Lao et al., 2022) and GraphGlow (Zhao et al., 2023) that are generative approaches utilizing variational inference technique.
@@
  • WSGNN (Lao et al., 2022) : https://github.com/Thinklab-SJTU/WSGNN
  • GraphGLOW (Zhao et al., 2023) : https://github.com/WtaoZhao/GraphGLOW
  • AirGNN (Liu et al., 2021) : https://github.com/lxiaorui/AirGNN
  • ProGNN (Jin et al., 2020) : https://github.com/ChandlerBang/Pro-GNN
  • RSGNN (Dai et al., 2022) : https://github.com/EnyanDai/RSGNN
  • STABLE (Li et al., 2022a) : https://github.com/likuanppd/STABLE
  • EvenNet (Lei et al., 2022) : https://github.com/Leirunlin/EvenNet
  • NRGNN (Dai et al., 2022) : https://github.com/EnyanDai/NRGNN
  • RTGNN (Dai et al., 2022) : https://github.com/GhostQ99/RobustTrainingGNN
  • SG-GSR (In et al., 2024) : https://github.com/yeonjun-in/torch-SG-GSR
@@ E.4 Evaluation Protocol
We mainly compare the robustness of Da-Gnn and the baselines under both the synthetic and real-world feature-dependent graph-noise (DANG). More details of generating DANG is provided in Sec E.2. Additionally, we consider independent feature/structure/label noise, which are commonly considered in prior works in this research field (Liu et al., 2021; Dai et al., 2022; Li et al., 2022a; In et al., 2023; Qian et al., 2023). Specifically, for the feature noise (Liu et al., 2021), we sample a subset of nodes (i.e., 10%, 30%, and 50%) and randomly flip 0/1 value on each dimension of node features $\mathbf{X}_{i}$ from Bernoulli distribution with probability $p=\frac{1}{F}\sum_{i=1}^{F}\mathbf{X}_{i}$. For the structure noise, we adopt the random perturbation method that randomly injects non-connected node pairs into the graph (Li et al., 2022a). For the label noise, we generate uniform label noise following the existing works (Qian et al., 2023; Dai et al., 2021).
@@ E.5 Implementation Details
For Da-Gnn, the learning rate is tuned from {0.01, 0.005, 0.001, 0.0005}, and dropout rate and weight decay are fixed to 0.6 and 0.0005, respectively. In the inference of $Z_{A}$, we use a 2-layer GCN model with 64 hidden dimension as $\text{GCN}_{\phi_{1}}$ and the dimension of node embedding $d_{1}$ is fixed to 64. The $\gamma$ value in calculating $\gamma$-hop subgraph similarity is tuned from {0, 1} and $k$ in generating $k$-NN graph is tuned from {0, 10, 50, 100, 300}. In the inference of $Z_{Y}$, we use a 2-layer GCN model with 128 hidden dimension as $\text{GCN}_{\phi_{3}}$. In the inference of $\epsilon_{X}$, the hidden dimension size of $\epsilon_{X}$, i.e., $d_{2}$, is fixed to 16. In the inference of $\epsilon_{A}$, the early-learning phase is fixed to 30 epochs. In the implementation of the loss term $-\mathbb{E}_{Z_{A}\sim q_{\phi_{1}}}\mathbb{E}_{\epsilon\sim q_{\phi_{2}}}\left[\log(p_{\theta_{1}}(A|X,\epsilon,Z_{A}))\right]$, we tune the $\theta_{1}$ value from {0.1, 0.2, 0.3}. In the overall learning objective, i.e., Eqn 3, $\lambda_{1}$ is tuned from { 0.003, 0.03. 0.3, 3, 30 }, $\lambda_{2}$ is tuned from { 0.003, 0.03. 0.3 }, and $\lambda_{3}$ is fixed to 0.001. We report the details of hyperparameter settings in Table 5.
[…]

### v2

@@ B.1 Details of regularizing the inference of $Z_{A}$
We regularize the learned latent graph $\mathbf{\hat{A}}$ based on the prior knowledge that pairs of nodes with high $\gamma$-hop subgraph similarity are more likely to form assortative edges zhao2023graphglow ; consm ; rsgnn , thereby encouraging $\mathbf{\hat{A}}$ to predominantly include such edges.
@@ Appendix C Further Discussion on DANG
C.1 Statistical Analysis on Evidence of DANG
To provide empirical evidence of the DANG assumption in addition to intuition, we conduct a statistical analysis on a real-world news network, PolitiFact wu2023decor , where node features represent news content, node labels correspond to news topics or categories, and edges denote co-tweet relationships—that is, instances where the same user tweeted both pieces of news. The network includes both fake and benign news, with fake news regarded as feature noise induced by malicious user intent ($\epsilon_{X}$).
To investigate noise dependency patterns associated with the presence of fake news, We hypothesize that the presence of fake news (i.e., feature noise) leads to noisy graph structures and noisy node labels in news networks. Specifically, we assign a semantic topic to each news article as a node label using k-means clustering over BERT embeddings of the article content. For each node in the graph, we compute the Shannon entropy of the semantic topic distribution among its neighboring nodes. We then compare these entropy values between fake and benign news nodes.
Table 4: Statistics comparison between Fake news and Benign news.
Category Mean 25% 50% (Median) 75%
Fake news 1.330 1.402 1.465 1.494
Benign news 1.084 1.004 1.353 1.431
Mann–Whitney U test (p-value) 4.19e-25
In Table 4, descriptive statistics reveal that fake news nodes generally exhibit higher entropy than benign news nodes, suggesting that benign news tends to connect to semantically similar articles (homophilic), whereas fake news is more frequently connected to semantically dissimilar articles (heterophilic). This observation aligns with common user behavior: people typically share news related to their interests, whereas fake news is often propagated indiscriminately, regardless of topical relevance kim2024revisiting . Furthermore, a non-parametric statistical test (Mann–Whitney U test) confirms that the difference in entropy values between fake and benign news is statistically significant. These findings suggest that the presence of fake news (i.e., node feature noise) introduces noisy and heterophilic edges into the graph structure. Furthermore, model-based automated news topic prediction often performs poorly due to noise in both features and graph structures, ultimately resulting in incorrect label annotations.
In summary, these findings empirically support the noisy dependency scenario in real-world scenario where feature noise (i.e., fake news content) can propagate through the graph, generating noisy edges and noisy labels. This highlights the need for our work that explicitly model and mitigate such noise dependencies in real-world networks.
C.2 Intuitive Examples of DANG
@@
  • Biological Networks: In addition to the user-item graphs, DANG manifests in the domain of single-cell RNA-sequencing (scRNA-seq). Specifically, in this graph the primary resource is a cell-gene count matrix. A cell-cell graph is commonly employed for downstream tasks, where each cell and its corresponding gene expression are represented as a node and node feature, respectively, and the cell type is considered a node label. However, the node feature, representing gene expression derived from the cell-gene count matrix, often contains noise due to various reasons, such as the dropout phenomenon dropout and batch effect batcheffect . Since the cell-gene count matrix is the main resource for generating the cell-cell graph wang2021scgnn ; xiong2023scgcl ; yun2023scFP , such noise acts as a significant obstacle in designing an effective graph structure. Additionally, cell types are annotated using transcripted marker genes, which serve as distinctive features characterizing specific cell types. Noisy node features, therefore, can lead to the misprediction of cell types (node labels). This issue of noise in node features in the biological domain underscores the critical challenge in real-world scenarios.
C.3 Extension of DANG
@@ Appendix D Further Discussion on DA-GNN
While the implementation of DA-GNN draws inspiration from the spirit of VAE kingma2013auto and CausalNL yao2021instance , we address complex and unique challenges absent in kingma2013auto ; yao2021instance . Specifically, the incorporation of $A$ necessitates handling supplementary latent variables and causal relationships, such as $Z_{A}$, $\epsilon_{A}$, $\epsilon_{A}, AA$$X$, YYAA, $AZ_{A}$, each posing non-trivial obstacles beyond their straightforward extension.
  • While yao2021instance assumes that $\epsilon$ only causes $X$, DANG posits that $\epsilon$ also causes $A$, denoted as $A\epsilon_{A}$. Consequently, DANG requires a novel inference/regularization approach for $\epsilon_{A}$, which is not addressed in yao2021instance , presenting a distinctive technical challenge.
  • A simplistic uniform prior is employed to regularize the modeling of $Z_{Y}$ in yao2021instance . However, upon close examination of the relationship YYAA, we advocate for a novel regularization approach for $Z_{Y}$ based on the principle of homophily. This method cannot be elicited through a straightforward application of yao2021instance to the graph.
  • By incorporating $A$, $Z_{A}$, and their associated casualties, we address distinct technical challenges, specifically the inference/regularization of $Z_{A}$ and the generation of $A$, which cannot be accommodated by a mere extension of yao2021instance to the graph. In particular, we utilize graph structure learning to model $Z_{A}$, and frame the generation of $A$ as an edge prediction task, incorporating novel regularization techniques for both edge prediction and label. Moreover, we regularize $Z_{A}$ leveraging our novel prior knowledge to enhance the accuracy and scalability of inference.
We argue that these components are non-trivial to handle through a straightforward application of yao2021instance to the graph domain.
@@ E.1 Datasets
We evaluate DA-GNN and baselines on five existing datasets (i.e., Cora coraciteseer , Citeseer coraciteseer , Amazon Photo and Computers shchur2018pitfalls ), and ogbn-arxiv hu2020open and two newly introduced datasets (i.e., Amazon Auto and Amazon Garden) that are proposed in this work based on Amazon review data amazondata1 ; amazondata2 to mimic DANG caused by malicious fraudsters on e-commerce systems (Refer to Appendix E.2.2 for details). The statistics of the datasets are given in Table 5. These seven datasets can be found in these URLs:
@@ E.2.2 Real-world DANG
We have introduced and released two new graph benchmark datasets, i.e., Auto and Garden, that simulate real-world DANG scenarios on e-commerce systems. To construct these graphs, we utilized metadata and product review data from two categories, "Automotives" and "Patio, Lawn and Garden," obtained from Amazon product review data sources amazondata1 ; amazondata2 . Specifically, we generated a clean product-product graph where node features are represented using a bag-of-words technique applied to product reviews. The edges indicate co-purchase relationships between products that have been purchased by the same user, and the node labels correspond to product categories. We perform both node classification and link prediction tasks, which are equivalent to categorizing products and predicting co-purchase relationships, respectively.
@@ E.3 Baselines
We compare DA-GNN with a wide range of noise-robust GNN methods, which includes feature noise-robust GNNs (i.e., AirGNN airgnn ), structure-noise robust GNNs (i.e., ProGNN prognn , RSGNN rsgnn , STABLE stable and EvenNet lei2022evennet ), label noise-robust GNNs (i.e., NRGNN nrgnn and RTGNN rtgnn ), and multifaceted noise-robust GNNs (i.e., SG-GSR in2024self ). We also consider WSGNN lao2022variational and GraphGlow zhao2023graphglow that are generative approaches utilizing variational inference technique.
@@
  • WSGNN lao2022variational : https://github.com/Thinklab-SJTU/WSGNN
  • GraphGLOW zhao2023graphglow : https://github.com/WtaoZhao/GraphGLOW
  • AirGNN airgnn : https://github.com/lxiaorui/AirGNN
  • ProGNN prognn : https://github.com/ChandlerBang/Pro-GNN
  • RSGNN rsgnn : https://github.com/EnyanDai/RSGNN
  • STABLE stable : https://github.com/likuanppd/STABLE
  • EvenNet lei2022evennet : https://github.com/Leirunlin/EvenNet
  • NRGNN rsgnn : https://github.com/EnyanDai/NRGNN
  • RTGNN rsgnn : https://github.com/GhostQ99/RobustTrainingGNN
  • SG-GSR in2024self : https://github.com/yeonjun-in/torch-SG-GSR
@@ E.4 Evaluation Protocol
We mainly compare the robustness of DA-GNN and the baselines under both the synthetic and real-world feature-dependent graph-noise (DANG). More details of generating DANG is provided in Sec E.2. Additionally, we consider independent feature/structure/label noise, which are commonly considered in prior works in this research field airgnn ; rsgnn ; stable ; in2023similarity ; rtgnn . Specifically, for the feature noise airgnn , we sample a subset of nodes (i.e., 10%, 30%, and 50%) and randomly flip 0/1 value on each dimension of node features $\mathbf{X}_{i}$ from Bernoulli distribution with probability $p=\frac{1}{F}\sum_{i=1}^{F}\mathbf{X}_{i}$. For the structure noise, we adopt the random perturbation method that randomly injects non-connected node pairs into the graph stable . For the label noise, we generate uniform label noise following the existing works rtgnn ; nrgnn .
@@ E.5 Implementation Details
For DA-GNN, the learning rate is tuned from {0.01, 0.005, 0.001, 0.0005}, and dropout rate and weight decay are fixed to 0.6 and 0.0005, respectively. In the inference of $Z_{A}$, we use a 2-layer GCN model with 64 hidden dimension as $\text{GCN}_{\phi_{1}}$ and the dimension of node embedding $d_{1}$ is fixed to 64. The $\gamma$ value in calculating $\gamma$-hop subgraph similarity is tuned from {0, 1} and $k$ in generating $k$-NN graph is tuned from {0, 10, 50, 100, 300}. In the inference of $Z_{Y}$, we use a 2-layer GCN model with 128 hidden dimension as $\text{GCN}_{\phi_{3}}$. In the inference of $\epsilon_{X}$, the hidden dimension size of $\epsilon_{X}$, i.e., $d_{2}$, is fixed to 16. In the inference of $\epsilon_{A}$, the early-learning phase is fixed to 30 epochs. In the implementation of the loss term $\mathbb{E}_{Z_{A}\sim q_{\phi_{1}}}\mathbb{E}_{\epsilon\sim q_{\phi_{2}}}\left[\log(p_{\theta_{1}}(A|X,\epsilon,Z_{A}))\right]$, we tune the $\theta_{1}$ value from {0.1, 0.2, 0.3}. In the overall learning objective, i.e., Eqn 4, $\lambda_{1}$ is tuned from { 0.003, 0.03. 0.3, 3, 30 }, $\lambda_{2}$ is tuned from { 0.003, 0.03. 0.3 }, and $\lambda_{3}$ is fixed to 0.001. We report the details of hyperparameter settings in Table 6.
For all baselines, we follow the training instruction reported in their paper and official code. For AirGNN, we tune $\lambda\in$ {0.0, 0.2, 0.4, 0.6, 0.8} and set the others as mentioned in the paper for all datasets. For ProGNN, we use the training script reported in the offical code since there are no training guidance in the paper. For RSGNN, we tune $\alpha\in$ {0.003, 0.03, 0.3, 3, 30}, $\beta\in$ {0.01, 0.03, 0.1, 0.3, 1}, $n_{p}\in$ {0, 10, 100, 300, 400}, and learning rate $\in$ {0.01, 0.005, 0.001, 0.0005} for all datasets. For STABLE, We tune $t_{1}\in$ {0.0, 0.01, 0.02, 0.03, 0.04}, $t_{2}\in$ {0.1, 0.2, 0.3}, $k\in$ {1, 3, 5, 7, 11, 13}, and $\alpha\in$ {-0.5, -0.3, -0.1, 0.1, 0.3, 0.6} for all datasets. For EvenNet, we tune $\lambda\in$ {0.1, 0.2, 0.5, 0.9} for all datasets following the training script of the official code. For NRGNN, we tune $\alpha\in$ {0.001, 0.01, 0.1, 1, 10}, $\beta\in$ {0.001, 0.01, 0.1, 1, 10, 100}, and learning rate $\in$ {0.01 , 0.005, 0.001, 0.0005} for all datasets. For RTGNN, we tune $K\in$ {1, 10, 25, 50, 100}, $th_{pse}\in$ {0.7, 0.8, 0.9, 0.95}, $\alpha\in$ {0.03, 0.1, 0.3, 1}, and $\gamma\in$ {0.01, 0.1}, and learning rate $\in$ {0.01, 0.005, 0.001, 0.0005}. For WSGNN, we use the best hyperparameter setting reported in the paper since there are no training guidance in the paper. For GraphGLOW, we tune learning rate $\in$ {0.001, 0.005, 0.01, 0.05}, embedding size $d\in$ {16, 32, 64, 96}, pivot number $P\in$ {800, 1000, 1200, 1400}, $\lambda\in$ {0.1, 0.9}, $H\in$ {4, 6}, $E\in\{1,2,3\}$, $\alpha\in$ {0, 0.1, 0.15, 0.2, 0.25, 0.3}, and $\rho\in$ {0, 0.1, 0.15, 0.2, 0.25, 0.3}. For SG-GSR, we tune learning rate $\in$ {0.001, 0.005, 0.01, 0.05}, $\lambda_{\mathcal{E}}$ $\in$ {0.2, 0.5, 1, 2, 3, 4, 5}, $\lambda_{\text{sp}}$ and $\lambda_{\text{fs}}$ $\in$ {1.0, 0.9, 0.7, 0.5, 0.3}, and $\lambda_{\text{aug}}$ $\in$ {0.1, 0.3, 0.5, 0.7, 0.9}.
@@ Appendix F Additional Experimental Results
F.1 Complexity Analysis
Large Scale Graph. To demonstrate the scalability of DA-GNN, we consider a larger graph dataset, ogbn-arxiv hu2020open . Table 1 clearly illustrates that DA-GNN exhibits superior scalability and robustness in comparison to other baseline methods.
Table 7: Training time comparison on Cora dataset under DANG 50%.
Total (sec) 20.9 702.1 159.9 53.3 0.8 100.3 118.7 86.3 46.3
per epoch (sec) 0.04 1.77 0.16 - 0.004 0.20 0.18 0.11 0.09
F.2 Sensitivity Analysis
We analyze the sensitivity of the coefficient $\lambda_{1}$ and $\lambda_{2}$ in Eqn 4, $\theta$, and $\gamma$. To be specific, we increase $\lambda_{1}$ value from $\{0.0,0.003,0.03,0.3,3\}$, $\lambda_{2}$ value from $\{0.0,0.003,0.03,0.3\}$, $\theta$ from $\{0.1,0.2,0.3\}$, and $\gamma$ from $\{0,1\}$. We then evaluate the node classification accuracy of DA-GNN under DANG.
  • In Fig 8(a) and 9(a), we notice that DA-GNN consistently surpasses the state-of-the-art baseline, EvenNet, regardless of the $\lambda_{1}$ value, demonstrating the robustness of DA-GNN. Furthermore, we observe that the performance significantly drops when $\lambda_{1}=0$. This highlights the importance of modeling the causal relationship $A\leftarrow(X,\epsilon,Z_{A})$ for robustness under DANG, as $\lambda_{1}$ is directly related to the loss term $\mathcal{L}_{\text{edge-rec}}$, i.e., $\mathbb{E}_{Z_{A}}\mathbb{E}_{\epsilon}\left[\log(p_{\theta_{1}}(A|X,\epsilon,Z_{A}))\right]$.
  • In Fig 8(b) and 9(b), we observe that DA-GNN generally outperforms the sota baseline regardless of the value of $\lambda_{2}$, indicating the stability of DA-GNN. Moreover, we can see a performance decrease when $\lambda_{2}=0$. This observation suggests that the regularization on the inferred latent node label $Z_{Y}$ using the inferred latent structure $Z_{A}$ effectively handles the noisy labels. This conclusion is drawn from the fact that $\lambda_{2}$ is directly linked to the loss term $\mathcal{L}_{\text{hom}}$, i.e., $kl(q_{\phi_{3}}(Z_{Y}|X,A)||p(Z_{Y}))$.
  • In Fig 8(c) and 9(c), we analyze the hyperparameter sensitivity of $k$ and observe that $k$ plays a critical role and requires some tuning. To recap the role of $k$, we pre-define a proxy graph based on subgraph similarity, where each node connects to $k$ neighbors. We then compute $\hat{p}$ as the edge weights on this proxy graph, which corresponds to the regularization term minimizing $kl(q_{\phi_{1}}(Z_{A}|X,A)||p(Z_{A})$. Sensitivity to $k$ highlights the importance of accurately inferring the latent graph structure $Z_{A}$. This is expected, as using rich neighborhood information from $Z_{A}$ enables robust message passing, which helps mitigate noise in the observed graphs. We restrict the search to just five values: {0, 10, 50, 100, 300}. This narrow range consistently yielded effective performance across all seven datasets, suggesting that tuning $k$ is not overly burdensome.
  • In Fig 8(d) and 9(d), we observe that DA-GNN consistently outperforms the state-of-the-art baseline, EvenNet, across all values of $\theta$, demonstrating the robustness of the prediction regularization method in Eqn3.
  • In Fig 8(e) and 9(e), we observe that DA-GNN consistently surpasses the state-of-the-art baseline, EvenNet, across all values of $\gamma$, emphasizing the stability of regularizing the inferred $Z_{A}$ in modeling $q_{\phi_{1}}(Z_{A}|X,A)$.
Figure 9: Sensitivity analysis on $\lambda_{1}$, $\lambda_{2}$, $\theta$, and $\gamma$. We conduct the experiments on Computers dataset under DANG-30%
F.3 Robustness Evaluation under Variants of DANG
F.3.1 Variants of Synthetic DANG
In the generation process of our synthetic DANG, we have three variables: 1) the overall noise rate, 2) the amount of noise dependency ($X\rightarrow A$, $X\rightarrow Y$, $A\rightarrow Y$), and 3) the amount of independent structure noise ($\epsilon\rightarrow A$).
Table 8: Node classification results on DANG with increased noise dependency
Dataset Setting AirGNN RSGNN STABLE EvenNet NRGNN RTGNN SG-GSR DA-GNN
Cora DANG-10% 78.3±0.3 79.0±0.1 79.5±0.4 76.9±1.2 78.6±0.5 78.8±0.5 78.5±0.2 79.8±0.2
DANG-30% 57.6±0.5 67.9±0.6 65.2±1.4 55.8±1.3 63.8±0.9 66.1±0.6 56.9±0.7 67.0±0.3
DANG-50% 40.1±0.5 49.4±0.9 45.7±1.2 40.5±1.0 47.5±0.5 48.1±0.8 40.1±1.2 51.6±0.9
Citeseer DANG-10% 65.7±1.1 72.9±0.4 68.4±0.7 68.8±0.6 69.7±1.0 69.8±0.0 70.3±0.5 72.7±0.3
DANG-30% 57.2±0.9 63.3±0.6 57.2±0.1 57.2±0.5 59.6±0.7 60.1±0.7 62.0±0.9 64.9±0.6
DANG-50% 39.8±0.7 49.4±1.0 41.3±1.8 42.2±0.5 42.9±0.6 43.7±0.7 46.1±1.3 51.4±0.2
Table 9: Node classification results on DANG without noise dependency
Dataset Setting AirGNN RSGNN STABLE EvenNet NRGNN RTGNN SG-GSR DA-GNN
Cora DANG-10% 83.1±0.4 83.9±0.6 83.9±0.4 84.0±0.5 83.8±0.2 84.8±0.2 84.8±0.1 84.8±0.1
DANG-30% 77.7±0.5 79.9±0.4 76.8±0.6 74.8±0.6 77.8±0.8 80.0±0.1 80.1±0.1 80.1±0.3
DANG-50% 66.9±2.6 72.7±0.6 70.3±1.8 61.3±3.4 69.1±0.8 73.2±0.6 72.0±0.2 75.4±0.0
Citeseer DANG-10% 68.8±0.3 75.9±0.9 71.9±0.6 74.0±0.3 74.1±0.7 74.4±0.4 75.5±0.5 75.9±0.4
DANG-30% 63.6±0.2 70.7±0.5 67.3±0.2 67.9±0.3 69.7±0.3 69.4±0.6 72.0±0.4 71.5±0.3
DANG-50% 59.7±0.7 64.6±0.7 59.3±0.6 61.3±0.6 63.4±0.4 64.6±0.2 66.4±0.3 64.7±0.6
Table 10: Node classification results on DANG without independent structure noise
Dataset Setting AirGNN RSGNN STABLE EvenNet NRGNN RTGNN SG-GSR DA-GNN
Cora DANG-10% 81.1±0.5 81.1±0.6 83.1±0.7 81.4±0.3 82.1±0.3 82.6±0.2 82.5±0.1 83.9±0.3
DANG-30% 73.9±1.7 73.6±0.3 76.9±0.3 69.7±0.7 76.3±0.4 74.8±0.9 77.7±0.3 79.6±0.6
DANG-50% 64.6±2.3 60.3±1.3 66.4±0.6 51.6±0.5 64.4±1.0 62.8±0.6 69.5±1.0 72.1±0.4
Citeseer DANG-10% 68.3±0.6 71.8±0.7 72.4±0.9 72.8±0.1 73.1±0.3 73.7±0.2 74.5±0.4 74.7±0.1
DANG-30% 58.5±0.5 63.5±0.9 64.6±0.2 63.1±0.4 64.3±1.4 64.8±0.9 66.1±0.6 66.4±0.6
DANG-50% 54.3±0.2 55.9±0.3 58.1±1.0 51.2±2.1 56.7±0.2 56.6±0.9 59.3±0.6 60.3±1.2
Table 11: Node classification results on DANG with increased independent structure noise
Dataset Setting AirGNN RSGNN STABLE EvenNet NRGNN RTGNN SG-GSR DA-GNN
Cora DANG-10% 78.9±0.7 81.6±0.3 80.9±0.5 78.9±0.3 79.9±0.4 81.8±0.3 81.4±0.2 82.5±0.2
DANG-30% 66.1±1.8 70.6±0.9 72.0±0.8 61.0±0.9 72.2±0.6 70.1±0.6 72.4±0.2 75.4±0.4
DANG-50% 47.3±0.5 56.7±0.2 57.9±1.6 42.1±2.1 58.8±0.6 55.6±0.5 61.2±1.5 65.4±0.6
Citeseer DANG-10% 66.5±0.4 74.0±0.3 70.8±0.4 70.3±0.8 71.9±0.1 72.6±0.3 72.6±0.4 73.6±0.2
DANG-30% 58.0±0.2 63.8±0.6 62.3±1.8 60.1±0.3 61.6±0.9 63.2±0.5 63.7±0.8 65.1±0.5
DANG-50% 49.4±0.6 55.1±0.3 51.7±1.7 45.9±0.7 50.8±0.8 52.0±1.1 54.2±0.2 55.4±1.2
  • For the first variable, our experiments already addressed it by varying the noise rate from 0 to 50.
  • For the second variable, we conduct an additional analysis by substantially increasing or decreasing the degree of noise dependency. Specifically, we increase the number of structure noise edges caused by feature noise by approximately 4×, and similarly amplify the amount of label noise induced by both feature and structure noise by 4×. We also evaluate a setting where noise dependencies are completely removed—this corresponds to a scenario with independent feature and structure noise. As shown in Table 8 and Table 9, DA-GNN consistently outperforms all baselines on the strong presence of noise dependency, and shows competitive performance on the weak presence of noise dependency.
  • For the third variable, we perform an additional analysis by doubling the amount of independent structure noise. We also evaluate the case where no independent structure noise is present. As shown in Table 10 and Table 11, DA-GNN consistently outperforms all baselines across both settings.
These results demonstrate that DA-GNN consistently outperforms other baselines under varying degrees of DANG, highlighting its practical applicability across diverse real-world noise conditions.
F.3.2 Variants of Real-world DANG
We conduct an experiment where we independently double each of the following: (1) the number of fraudsters (i.e., nodes with noisy features) and (2) the activeness of fraudsters (i.e., the amount of structure noise they introduce) in our real-world DANG generation process. As a result, label noise also increases accordingly, in proportion to the amount of generated feature and structure noise.
Table 12: Node classification results on variants of real-world DANG
Dataset Setting AirGNN RSGNN STABLE EvenNet NRGNN RTGNN DA-GNN
Auto DANG w/ doubled # frauds 54.6±1.5 53.4±0.7 55.4±0.1 56.5±0.6 55.9±1.5 54.3±2.7 60.1±0.7
DANG w/ doubled structure noise 56.9±0.7 50.9±0.6 58.1±2.0 53.6±2.2 55.8±1.3 56.78±0.8 55.6±1.0
Garden DANG w/ doubled # fraudsters 57.1±1.3 65.0±0.5 69.8±2.3 69.3±1.8 70.8±0.7 70.6±0.9 71.9±0.6
DANG w/ doubled structure noise 69.9±2.8 69.4±1.3 72.0±0.5 72.4±0.9 71.0±2.4 75.3±0.4 74.4±0.2
As shown in Table 12, DA-GNN demonstrates competitive performance and, in many cases, outperforms other baselines under these intensified noise conditions.